OpenStudy (anonymous):

An open rectangular box with square base is to be made from 48 ft2 of material. What dimensions will result in a box with the largest possible volume? Please show work, I'm lost! THanks

OpenStudy (zzr0ck3r):

base is square with length = width = x, and height = y volume = x^2*y minimize f(x,y) = x^2*y subject to x^2+4xy = 48 (area of square base and 4 "walls" with area w*h) y = (48-x^2)/4x f(x,y) = f(x) = x^2((48-x^2)/4x) = x((48-x^2)/4 = (48x-x^3)/4 = 12x - (1/4)x^3 f'(x) = 0 --> 12 - (3/4)x^2 = 0 (3/4)x^2 = 12 x^2 = 16 x= +- 4 -4 does not make sense so x = 4 so x^2+4xy = 48 ---> 16+16y = 48 --> y = 2 so max volume is 4*4*2 = 32ft^3

OpenStudy (zzr0ck3r):

@skdavis93 let me know if you have probs.